OpenStudy (anonymous):

4 + 7(x - 12) = -38 x=

OpenStudy (abb0t):

distribute the 7 inside the parenthesis. then combine like terms.

OpenStudy (anonymous):

4+7x-84=-38

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

Yes, not combine like terms. You can add 4 and (-84)

OpenStudy (abb0t):

Then subtract on the left. Notice that you're isolating 7.

OpenStudy (anonymous):

7x+80=-38

OpenStudy (anonymous):

you make mistake at somewhere, I don't know where, but my answer is not the same with yours.

OpenStudy (anonymous):

im stuck @abb0t

OpenStudy (anonymous):

@Hoa im stuck

OpenStudy (anonymous):

let me show you mine, is it ok, or wait for abbot?

OpenStudy (anonymous):

go right ahead @Hoa

OpenStudy (anonymous):

4+7(x-12)=-38 ----> 7(x-12)= -38-4=-42 -----> x -12 = -42/7 = -6 ---> x = -6+12= 6 that's it

OpenStudy (anonymous):

THANKS @Hoa

OpenStudy (anonymous):

you are very welcome.